Is there any difference between light waves and the matter waves?

Grade:8

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
a) Light waves are relatively long; while matter waves are exceedingly short above atomic mass size. Light waves are born strictly from the kinetic energy of the photons. Matter waves come from both the mass, m, and the kinetic energy of the matter.
b) DB started with E = Mc^2 = hf = hc/L; so that L = h/M is the DB wave. Note that hf = E = Pc is the kinetic energy of photons with momentum P = hf/c. So both the DB and light waves have their origins in the particle energies.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty